# CheMixHub: Datasets and Benchmarks for Chemical Mixture Property Prediction

### v1

@@ Abstract
Developing improved predictive models for multi-molecular systems is crucial, as nearly every chemical product used results from a mixture of chemicals. While being a vital part of the industry pipeline, the chemical mixture space remains relatively unexplored by the Machine Learning community. In this paper, we introduce CheMixHub, a holistic benchmark for molecular mixtures, covering a corpus of 11 chemical mixtures property prediction tasks, from drug delivery formulations to battery electrolytes, totalling approximately 500k data points gathered and curated from 7 publicly available datasets. CheMixHub introduces various data splitting techniques to assess context-specific generalization and model robustness, providing a foundation for the development of predictive models for chemical mixture properties. Furthermore, we map out the modelling space of deep learning models for chemical mixtures, establishing initial benchmarks for the community. This dataset has the potential to accelerate chemical mixture development, encompassing reformulation, optimization, and discovery. The dataset and code for the benchmarks can be found at: https://github.com/chemcognition-lab/chemixhub
@@ 1 Introduction
Mixtures of molecules are integral to our daily experiences: from the perfumes we smell [58] to the remedies we take [3, 74]. Understanding the interactions and behaviors of molecular mixtures is therefore essential for advancements in biochemistry [17], drug discovery [3] and environmental science [24]. Mixtures are particularly appealing because they offer greater flexibility in tailoring properties that specific application needs. By adjusting composition, substituting components, or introducing new ones, it is possible to fine-tune characteristics such as viscosity [6], volatility [19], stability [36], and conductivity [32]—features that are highly task-dependent. However, discovering new chemical mixtures, optimizing or reformulating them requires thorough characterization, which is time- and resource-intensive due to the exponentially growing number of candidate combinations. The mixture search space is vastly larger than that of single-component systems, making exhaustive experimental exploration impractical.
While Machine Learning (ML) has emerged as a powerful tool for accelerating the characterization and discovery of new materials [49], it faces unique challenges in the context of mixtures. On one hand, mixtures properties are highly correlated with strength of intermolecular interactions that cannot be inferred directly from the behavior of individual components. This causes quantitative structure–property relationship (QSPR) modeling remain underexplored for multi-component systems compared to the substantial progress achieved for single-component systems [46]. On the other hand, the scarcity of publicly available datasets further limits progress in this area. While mono-molecular systems have benefited from well-established, community-driven datasets and benchmarks—such as MoleculeNet [68] and the Therapeutics Data Commons [25]—no centralized or standardized database currently exists for multi-component molecular systems.
@@ 2 Related works
Various open-access and commercial sources offer experimental and computational mixture data, though these predominantly cover binary and ternary systems, with complex multi-component mixtures underrepresented (see Figure 2) [59, 44, 73, 34]. For instance, the open-source NIST ILThermo database [27]provides temperature-dependent transport and thermophysical properties for ionic liquid mixtures. Commercial platforms like DETHERM [67] and the Dortmund Data Bank [40] offer thermophysical data but are not publicly accessible. Despite these resources, substantial mixture data remains scattered throughout the literature, highlighting the need for unified datasets to enable robust mixture behavior modeling [80, 7, 14] and drive progress.
@@
Predicting mixture properties from a set of components requires models that respect key inductive biases, notably permutation invariance [69] [23]. Many permutation-invariant set functions follow a common blueprint: a series of permutation-equivariant operations (e.g., element-wise MLPs or self-attention layers) followed by a permutation-invariant aggregation (e.g., summation or pooling) [8]. The pioneering DeepSets architecture [72, 41] exemplifies this, using an element-wise MLP and sum aggregation, followed by further non-linear processing ( sum-decomposition [51, 56, 47, 64, 38, 82]). SetTransformer [33, 13] uses self-attention to model pairwise interactions with attention-based aggregation. Janossy pooling [39] offers another approach, explicitly modeling invariance and capturing higher-order interactions by averaging outputs over multiple input permutations [69, 78].
@@
ML is increasingly applied to chemical mixtures, primarily for property prediction, with emerging work on optimization and discovery [81, 53]. Common strategies involve aggregating molecular-level chemo-informatic features, graph neural networks (GNNs) embeddings [54, 75, 58, 80, 6] or large pre-trained chemical language models (CLMs) [55, 45]—using a DeepSets-like architecture. Alternatively, mixture representations are formed by weighted combinations of individual component descriptors [54, 55, 45], or learned by tree-based models from these descriptors [3]. A separate line of research explores attention-based architectures [75, 58]. Some work explicitly models pairwise interaction terms between mixture components, offering a more physically grounded and expressive representation [80]. These mixture embeddings are then fed to predictive models like neural networks [6], gradient boosting machines [3], or Gaussian processes [53].
3 Datasets
3.1 Dataset overview
@@
We curated 11 regression tasks from 7 published datasets across the chemical mixture literature, which are summarized in Table 1. The tasks were selected based on application domain and prior use as baselines in ML studies. To ensure their accessibility and standardization for ML applications, we provide a Croissant [2] file detailing their metadata, structure, and semantics based on schema.org. The dataset licenses are listed in Section A.1.
@@
Table 1: CheMixHub tasks summary. T indicates Temperature dependency. Mole Fractions indicates if mole fractions of the mixture components are available. Arrhenius relationship indicates if the task can be modeled using the Arrhenius equation. Exp. indicates if the data was obtained from wet-lab experiments or simulations.
Tasks Units Datapoints Max # Components # Unique Mixtures # Unique Molecules Mixture Context Mole Fractions Arrhenius Relationship Exp.
Miscible solvents $\rho$ g/m3 30,142 5 19,238 81
$\Delta H_{mix}$ kJ/mol 30,142 5 19,238 81
$\Delta H_{vap}$ kcal/mol 30,142 5 19,238 81
IlThermo $\ln(\kappa)$ S/m 40,904 3 14,438 479 T
[…]
NIST $\ln(\eta_{source})$ cP 239,201 2 84,133 1648 T
$\ln(\eta_{mixprop})$ cP 34,374 2 4566 1397 T
Drug solubility $\ln(S)$ g/100g 27,166 3 3259 169 T
Solid Polymer Electrolyte $\ln(\kappa)$ S/m 11,350 5 1749 402 T
Olfactory similarity 865 43 743 201
Motor Octane Number (MON) 684 121 352 419
3.2 Tasks
Miscible Solvents (3 tasks)
Homogeneous solutions are important in a variety of material science applications such as battery electrolytes, chemical reactivity, and consumer packaged goods. The Miscible Solvents dataset is a set of three tasks centered around miscible solvent properties, originally generated by Chew et al. using molecular dynamics (MD) simulations for 19,238 unique mixtures [14].
  • Density ($\rho$): $\rho$ measures how tightly packed the molecules are for a mixture. In industrial applications, density is important as it dictates the final weight and polarity of the product.
  • Heat of vaporization ($\Delta H_{vap}$): $\Delta H_{vap}$ is the amount of heat needed to convert some fraction of liquid into vapor. While experimentally measuring $\Delta H_{vap}$ for mixtures is challenging, it effectively measures the cohesion energy of a liquid and has been previously observed to correlate with temperature-dependent viscosity [15].
  • Enthalpy of mixing ($\Delta H_{m}$): $\Delta H_{m}$ is a fundamental thermodynamic property of liquid mixtures that measures the energy released or absorbed upon the mixing of pure components into a single phase in equilibrium. It is important for process design that dictates properties, such as solubility and phase stability.
IlThermo (2 tasks)
Ionic liquids (ILs) are salts composed of organic cations and organic or inorganic anions that remain liquid at temperatures below 100 °C [1, 61, 79]. ILThermo is a web-based database that provides extensive information on over 50 chemical and physical properties of pure ILs, as well as their binary and ternary mixtures with various solvents [28]. For the scope of this paper, we selected two property prediction tasks from IlThermo; however, we have open-sourced our curation code to facilitate the addition of further tasks in the future. Details of the curation process are provided in Section A.7, and the selected tasks are summarized below:
  • Ionic conductivity ($\kappa$): Higher $\kappa$ makes ILs attractive for use as electrolytes in energy storage and other electrochemical applications [37]. However, ionic conductivity of ionic liquid mixtures is a complex phenomenon and is influenced by multiple factors such as size and charge on the ions, polarity and dielectric strength of the solvent, viscosity, hydrogen bonding strength, ion association, etc [57]. To facilitate data-driven approaches, IlThermo dataset includes 40,904 $\kappa$ data points curated from literature, covering 14,438 unique mixtures composed of 479 molecules.
  • Viscosity ($\eta$): Modeling the viscosity of ILs is particularly challenging, as their viscosity can be orders of magnitude higher than those of conventional solvents [43]. This complexity arises from the coexistence of multiple interaction types—ionic, dispersion, dipole-dipole, and induced dipole interactions—that are more pronounced compared to typical organic solvents. The ILThermo dataset provides 75,992 viscosity ($\eta$) data points curated from the literature, encompassing 15,878 unique mixtures formed from 699 distinct molecules.
NIST viscosity (2 task)
Dynamic viscosity is a key design objective for modern process engineering and products. However, modeling the viscosity of liquid mixtures presents significant challenges due to the complex molecular interactions and the potential for nonmonotonic behavior [50]. NIST Thermodynamics Research Center (TRC) data archival system provides one of the most comprehensive datasets containing 239,201 dynamic viscosity datapoints of binary liquid mixtures [27]. Bilodeau et al. proposed a smaller version of the dataset by applying two key preprocessing steps: (1) removing data entries with SMILES strings containing multiple, non-covalently bonded fragments, and (2) excluding entries where either molecule was predicted to be a gas or solid in its pure form. These steps reduced the dataset to 34,374 data points[6].
Drug solubility (1 task)
The drug solubility in mixture of solvents is a critical factor that influences various stages of the pharmaceutical development pipeline, from drug discovery, drug analysis to formulation design. It allows greater flexibility through adjusting solvent combinations and ratios enabling solubility to be tailored to meet specific needs and to co-dissolve other necessary materials. The drug solubility dataset originally curated by Bao et al. from literature includes 27,166 data points [3]
Solid polymer electrolytes conductivity (1 task)
Solid polymer electrolytes (SPEs), proposed as potential replacements for conventional liquid organic electrolytes in batteries, have been engineered to offer improved electrochemical stability and reduced flammability. However, their practical use is limited by inherently low ionic conductivity. To support data-driven research on this issue, Bradford et al. compiled a dataset from the literature comprising 11,350 ionic conductivity measurements across more than 1,700 unique electrolyte formulations. Each formulation is uniquely defined by the polymer, salt, salt concentration, polymer molecular weight, and any additives present [7].
Motor Octane Number (1 task)
Kuzhagaliyeva et al. compiled a database containing 684 data points for 352 unique single hydrocarbons and mixtures, reporting experimentally measured motor octane numbers (MON) from various literature sources. MON is a combustion-related property commonly used to assess a fuel’s resistance to knocking. The dataset is categorized into three subpopulations: pure components, blends with 10 or fewer components (mostly surrogates), and complex fuels containing more than 10 components [31].
Olfactory similarity (1 task)
The dataset was originally compiled from previous publications [66, 52, 9] by Tom et al. [58]. Data for each of these publications was obtained from pyrfume [11] and consists of 865 pairwise mixture comparisons. Each pair is assigned a continuous perceptual similarity score ranging from 0 (completely similar) to 1 (completely different). This final score represents an average of similarity ratings obtained from human participants across different experimental paradigms.
@@ 3.3 Curation pipeline
  • Number of components While our focus is on multi-component systems, we preserve the single-component data points in the datasets with the exception of IlThermo. We leave the choice to the user to filter out single-component data points in CheMixHub.
@@
  • Data scales Due to great variations in experimental value ranges, it is common to apply logarithmic scaling to conductivity, solubility and viscosity properties [6, 7, 3]. We follow this principle and apply it to these types of properties.

### v2

@@ Abstract
Developing improved predictive models for multi-molecular systems is crucial, as nearly every chemical product used results from a mixture of chemicals. While being a vital part of the industry pipeline, the chemical mixture space remains relatively unexplored by the Machine Learning (ML) community. In this paper, we introduce CheMixHub, a holistic benchmark for molecular mixtures spanning a corpus of 11 chemical mixtures property prediction tasks. With applications ranging from drug delivery formulations to battery electrolytes, CheMixHub currently totals approximately 500k data points gathered and curated from 7 publicly available datasets. We devise various data splitting techniques to assess context-specific generalization and model robustness, providing a foundation for the development of predictive models for chemical mixture properties. Furthermore, we map out the modelling space of deep learning models for chemical mixtures, establishing initial benchmarks for the community. This dataset has the potential to accelerate chemical mixture development, encompassing reformulation, optimization, and discovery. The dataset and code for the benchmarks can be found at: https://github.com/chemcognition-lab/chemixhub
@@ 1 Introduction
Mixtures of molecules are integral to our daily experiences: from the perfumes we smell [tom2025molecules] to the remedies we take [bao2024towards, zaslavsky2023dataset]. Understanding the interactions and behaviors of molecular mixtures is therefore essential for advancements in biochemistry [derby2022understanding], drug discovery [bao2024towards] and environmental science [heys2016risk]. Mixtures are particularly appealing because they offer greater flexibility in tailoring properties that specific application needs. By adjusting composition, substituting components, or introducing new ones, it is possible to fine-tune characteristics such as viscosity [bilodeau2023machine], volatility [evjen2019viscosity], stability [liu2023experimental], and conductivity [latini1996alternative]—features that are highly task-dependent. However, discovering new chemical mixtures, optimizing or reformulating them requires thorough characterization, which is time- and resource-intensive due to the exponentially growing number of candidate combinations. The mixture search space is vastly larger than that of single-component systems, making exhaustive experimental exploration impractical.
While ML has emerged as a powerful tool for accelerating the characterization and discovery of new materials [rajabi2025adaptive], it faces unique challenges in the context of mixtures. On one hand, mixtures properties are highly correlated with strength of intermolecular interactions that cannot be inferred directly from the behavior of individual components. This causes quantitative structure–property relationship (QSPR) modeling remain underexplored for multi-component systems compared to the substantial progress achieved for single-component systems [pyzer2022accelerating]. On the other hand, the scarcity of publicly available datasets further limits progress in this area. While mono-molecular systems have benefited from well-established, community-driven datasets and benchmarks—such as MoleculeNet [wu2018moleculenet] and the Therapeutics Data Commons [Huang2021tdc]—no centralized or standardized database currently exists for multi-component molecular systems.
@@ 2 Related works
Various open-access and commercial sources offer experimental and computational mixture data, though these predominantly cover binary and ternary systems, with complex multi-component mixtures underrepresented (see Figure 2) [uceda2025experimental, podgorsek2016mixing, zang2025excess, li2025thermophysical]. For instance, the open-source NIST ILThermo database [kazakov2012nist]provides temperature-dependent transport and thermophysical properties for ionic liquid mixtures. Commercial platforms like DETHERM [westhaus1999detherm] and the Dortmund Data Bank [onken1989dortmund] offer thermophysical data but are not publicly accessible. Despite these resources, substantial mixture data remains scattered throughout the literature, highlighting the need for unified datasets to enable robust mixture behavior modeling [zhu2024differentiable, bradford2023chemistry, chew2025leveraging] and drive progress.
@@
Predicting mixture properties from a set of components requires models that respect key inductive biases, notably permutation invariance [xie2025advances] [goyal2022inductive]. Many permutation-invariant set functions follow a common blueprint: a series of permutation-equivariant operations (e.g., element-wise MLPs or self-attention layers) followed by a permutation-invariant aggregation (e.g., summation or pooling) [2104.13478]. The pioneering DeepSets architecture [zaheer2017deep, ou2022learning] exemplifies this, using an element-wise MLP and sum aggregation, followed by further non-linear processing ( sum-decomposition [ravanbakhsh2016deep, taraday2024sequential, qi2017pointnet, wagstaff2019limitations, maron2020learning, zweig2022exponential]). SetTransformer [lee2019set, chen2024stacking] uses self-attention to model pairwise interactions with attention-based aggregation. Janossy pooling [murphy2018janossy] offers another approach, explicitly modeling invariance and capturing higher-order interactions by averaging outputs over multiple input permutations [xie2025advances, zhang2019fspool].
@@
ML is increasingly applied to chemical mixtures, primarily for property prediction, with emerging work on optimization and discovery [zohair2025chemical, ruza2025autonomous]. Common strategies involve aggregating molecular-level chemo-informatic features, graph neural networks (GNNs) embeddings [sharma2023formulation, zhang2024learning, tom2025molecules, zhu2024differentiable, bilodeau2023machine] or large pre-trained chemical language models (CLMs) [soares2023capturing, priyadarsini2024improving]—using a DeepSets-like architecture. Alternatively, mixture representations are formed by weighted combinations of individual component descriptors [sharma2023formulation, soares2023capturing, priyadarsini2024improving], or learned by tree-based models from these descriptors [bao2024towards]. A separate line of research explores attention-based architectures [zhang2024learning, tom2025molecules]. Some work explicitly models pairwise interaction terms between mixture components, offering a more physically grounded and expressive representation [zhu2024differentiable]. These mixture embeddings are then fed to predictive models like neural networks [bilodeau2023machine], gradient boosting machines [bao2024towards], or Gaussian processes [ruza2025autonomous].
3 Dataset
3.1 Corpus overview
@@
We curated 11 regression tasks from 7 published datasets across the chemical mixture literature, which are summarized in Table 1. The tasks were selected based on application domain and prior use as baselines in ML studies. To ensure their accessibility and standardization for ML applications, we provide a Croissant [10.1145/3650203.3663326] file detailing their metadata, structure, and semantics based on schema.org. The dataset licenses are listed in Appendix A.1. Additional statistics on the molecules found in each task are provided in Appendix A.2.
Table 1: CheMixHub tasks summary. T indicates temperature dependency. Mole Fractions indicates availability of mole fraction information. Arrhenius relationship indicates if the target property can be modeled using the Arrhenius equation. Exp. indicates if the data was obtained from wet-lab experiments or simulations.
Dataset Tasks
Units
Datapoints
Max #
Components
# Unique
Mixtures
# Unique
Molecules
Mixture
Context
Mole
Fractions
Arrhenius
Relationship
Exp.
Miscible Solvents $\rho$ g/m3 30,142 5 19,238 81
$\Delta H_{\mathrm{mix}}$ kJ/mol 30,142 5 19,238 81
$\Delta H_{\mathrm{vap}}$ kcal/mol 30,142 5 19,238 81
IlThermo
(Ionic Liquids)
$\ln(\kappa)$ S/m 40,904 3 14,438 479 T
[…]
NIST Viscosity
(Liquid mixtures)
$\ln(\eta_{\mathrm{NIST-full}})$ cP 239,201 2 84,133 1648 T
$\ln(\eta_{\mathrm{NIST}})$ cP 34,374 2 4566 1397 T
Drug solubility $\ln(S)$ g/100g 27,166 3 3259 169 T
Solid Polymer Electrolytes $\ln(\kappa)$ S/m 11,350 5 1749 402 T
Olfactory mixtures
Perceptual similarity
865 43 743 201
Fuel mixtures MON 684 121 352 419
3.2 Datasets & Tasks
Miscible Solvents (3 tasks)
Homogeneous solutions are important in a variety of material science applications such as battery electrolytes, chemical reactivity, and consumer packaged goods. The Miscible Solvents dataset provides a set of three tasks centered around miscible solvent properties, originally generated by Chew et al. using molecular dynamics (MD) simulations for 19,238 unique mixtures [chew2025leveraging].
  • Density ($\rho$): $\rho$ measures how tightly packed the molecules are in a mixture. In industrial applications, density is important as it dictates the final weight and polarity of the product.
  • Heat of vaporization ($\Delta H_{\mathrm{vap}}$): $\Delta H_{\mathrm{vap}}$ is the amount of heat needed to convert some fraction of liquid into vapor. While experimentally measuring $\Delta H_{\mathrm{vap}}$ for mixtures is challenging, it effectively measures the cohesion energy of a liquid and has been previously observed to correlate with temperature-dependent viscosity [chew2024advancing].
  • Enthalpy of mixing ($\Delta H_{\mathrm{mix}}$): $\Delta H_{\mathrm{mix}}$ is a fundamental thermodynamic property of liquid mixtures that measures the energy released or absorbed upon the mixing of pure components into a single phase in equilibrium. It is important for process design that dictates properties, such as solubility and phase stability.
IlThermo (2 tasks)
Ionic liquids (ILs) are salts composed of organic cations and organic or inorganic anions that remain liquid at temperatures below 100 °C [aghaie2018systematic, vekariya2017review, zhou2017ionic]. ILThermo is a web-based database that provides extensive information on over 50 chemical and physical properties of pure ILs, as well as their binary and ternary mixtures with various solvents [kazakov2013ionic]. For the scope of this paper, we selected two property prediction tasks from IlThermo; however, we have open-sourced our curation code to facilitate the addition of further tasks in the future. Details of the curation process are provided in Section A.3, and the selected tasks are summarized below:
  • Ionic conductivity ($\kappa$): Higher $\kappa$ makes ILs attractive for use as electrolytes in energy storage and other electrochemical applications [liu2023locally]. However, ionic conductivity of ionic liquid mixtures is a complex phenomenon and is influenced by multiple factors such as size and charge on the ions, polarity and dielectric strength of the solvent, viscosity, hydrogen bonding strength, ion association, etc [thorat2025identifying]. To facilitate data-driven approaches, IlThermo dataset includes 40,904 $\kappa$ data points curated from literature, covering 14,438 unique mixtures composed of 479 distinct molecules.
  • Viscosity ($\eta$): Modeling the viscosity of ILs is particularly challenging, as their viscosity can be orders of magnitude higher than those of conventional solvents [philippi2022pressing]. This complexity arises from the coexistence of multiple interaction types—ionic, dispersion, dipole-dipole, and induced dipole interactions—that are more pronounced compared to typical organic solvents. The ILThermo dataset provides 75,992 viscosity ($\eta$) data points curated from the literature, encompassing 15,878 unique mixtures formed from 699 distinct molecules.
NIST viscosity (2 task)
Dynamic viscosity is a key design objective for modern process engineering and products. However, modeling the viscosity of liquid mixtures presents significant challenges due to the complex molecular interactions and the potential for nonmonotonic behavior [ramirez2024viscosity]. NIST Thermodynamics Research Center (TRC) data archival system provides one of the most comprehensive datasets containing 239,201 dynamic viscosity datapoints of binary liquid mixtures [kazakov2012nist]. Bilodeau et al. proposed a smaller version of the dataset by applying two key preprocessing steps: (1) removing data entries with SMILES strings containing multiple, non-covalently bonded fragments, and (2) excluding entries where either molecule was predicted to be a gas or solid in its pure form. These steps reduced the dataset to 34,374 data points[bilodeau2023machine]. We include both version of the dataset and refer to each as NIST-full and NIST, respectively.
Drug solubility (1 task)
The drug solubility in mixture of solvents is a critical factor that influences various stages of the pharmaceutical development pipeline, from drug discovery, drug analysis to formulation design. It allows greater flexibility through adjusting solvent combinations and ratios enabling solubility to be tailored to meet specific needs and to co-dissolve other necessary materials. The dataset was originally curated by Bao et al. from literature and includes 27,166 data points [bao2024towards]
Solid polymer electrolytes ionic conductivity (1 task)
Solid polymer electrolytes (SPEs), proposed as potential replacements for conventional liquid organic electrolytes in batteries, have been engineered to offer improved electrochemical stability and reduced flammability. However, their practical use is limited by inherently low ionic conductivity. To support research on this issue, Bradford et al. compiled a dataset from the literature comprising 11,350 ionic conductivity measurements across more than 1,700 unique electrolyte formulations. Each formulation is uniquely defined by the polymer, salt, salt concentration, polymer molecular weight, and any additives present [bradford2023chemistry].
Fuel mixture Motor Octane Number (1 task)
Kuzhagaliyeva et al. compiled a database containing 684 data points for 352 unique single hydrocarbons and mixtures, reporting experimentally measured motor octane numbers (MON) from various literature sources. The MON is a combustion-related property commonly used to assess a fuel’s resistance to knocking. The dataset is categorized into three subpopulations: pure components, blends with 10 or fewer components (mostly surrogates), and complex fuels containing more than 10 components [kuzhagaliyeva2022artificial].
Olfactory mixture perceptual similarity (1 task)
Predicting the perceptual similarity of olfactory mixture contributes to olfaction digitization efforts [lee2023principal] and also enables mixture reformulation. The dataset was originally compiled from previous publications [weiss2012perceptual, ravia2020measure, bushdid2014humans] by Tom et al. [tom2025molecules]. Data for each of these publications was obtained from pyrfume [castro2022pyrfume] and consists of 865 pairwise mixture comparisons. Each pair is assigned a continuous perceptual similarity score ranging from 0 (completely similar) to 1 (completely different). This final score represents an average of similarity ratings obtained from human participants across different experimental paradigms.
@@ 3.3 Curation pipeline
  • Handling of chemical 3D geometry In the datasets we consider, the majority of molecules possess fewer than five rotatable bonds (see Appendix A.2). This limits the expected benefit of conformer ensemble approaches in our context [zhu2023learning]. We leave the study of the impact of 3D conformations information on property prediction of mixtures of highly flexible molecules for future work.
  • Number of components While our focus is on multi-component systems, we preserve the single-component data points in the datasets with the exception of IlThermo (see Figure 2). We leave the choice to the user to filter out single-component data points in CheMixHub.
@@
  • Data scales Due to great variations in experimental value ranges, it is common to apply logarithmic scaling to conductivity, solubility and viscosity properties [bilodeau2023machine, bradford2023chemistry, bao2024towards]. We follow this principle and apply it to these types of properties.